# Casimir Arc–Plate Geometry: Computational Analysis of Thickness Constraints for Gold and Silver Nanomembranes in MEMS Applications

### v1

@@ 2.4 Procedure
In classical plate and shell theory, curvature reversal corresponds to a buckling-type instability whose critical condition depends on boundary constraints and in-plane stresses. A full buckling analysis under Casimir loading would require solving nonlinear equilibrium equations and identifying bifurcation points for curved beams or shells, typically involving eigenvalue problems and numerical methods [28]. Such an analysis is beyond the scope of our study and is left for future investigation.
@@ 2.4.1 Calculating Casimir Energy
Fosco’s correction was implemented by applying a derivative expansion to the first two terms as described in equation 13 where $U_{DE}$ is the Casimir energy including the NTLO correction, $\hbar$ is the reduced Planck constant, $c$ is the speed of light in vacuum, $x_{||}$ denotes the coordinates parallel to the reference surface over which the integration is performed, $\psi(d)$ is the local separation, $d$, between the curved surface and the opposing plate (equation 14), and $(\partial\psi)^{2}$ represents the sum of the squared partial derivatives of $\psi$ with respect to the parallel coordinates, i.e., $(\partial\psi)^{2}=(\partial\psi/\partial x)^{2}+(\partial\psi/\partial y)^{2}$ [8].
The geometric function describes the arc, as expressed in equation 14. Restrictions were imposed, namely [3<y<3][-3<y<3], to prevent contact between the two fixed surfaces. As described in [8], the first term in the derivative expansion corresponds to PFA. Thus, if one approximates the second term as a small perturbation parameter, $\epsilon$, by perturbation theory, $U_{DE}$ reduces the expression to the PFA.
NTLO corrections serve to better approximate the geometry of the arc due to its geometrical properties, specifically the tendency of a slope which can change more rapidly where local distances change nonlinearly. Experimental proof of the accuracy of PFA with NTLO corrections is left for future developments.
@@ 2.4.2 Calculating Bending Energy
The bending energy, $U_{bending}$, is obtained using equation 15, where $u$ is the bending strain energy and $A$ is the area [29]. Because we obtain a numerical value for $u$, the equation simplifies to equation 16, where $A$ simplifies to $L$, the length of the arc. To evaluate the arc length, $L$ is defined in equation 17, where $y_{min}=-3\mu m$ and $y_{max}=3\mu m$. The length of the arc was $L=6\mu m$.
@@
Table 1 shows the values of $E$, Young’s modulus, and $\upsilon$, Poisson’s ratio, for thin gold and thin silver films [30], [31].
@@ 2.5 Algorithm for Critical Thickness Evaluation
The procedure for evaluating the critical thickness can be summarized in algorithm 1.
@@
7:Apply derivative expansion correction [8]:
@@ 3.1 Estimating the maximum thickness using PFA with NTLO corrections
By setting the perturbation parameter, $\epsilon=0.1$, the results of PFA (without NTLO corrections) were obtained. The results for the PFA approximation suggest that the maximum thickness of both the silver and gold nanomembranes is smaller than the estimation with the NTLO corrections. In fact, the calculations without NTLO corrections suggest that the thickness would be below the atomic thickness of both gold and silver (0.409 nm for a single atomic layer of silver, and 0.47 nm for a two-atom-thick gold nanosheet). Such extremely low values are clearly beyond the physically meaningful range for continuous membranes and indicate that higher-order corrections are necessary.
@@
Including higher-order corrections to the PFA contributes to a more realistic estimate of the results, as dictated by the estimate of the fractional deviation, which is equal for both the silver and gold calculations. The fractional deviation $\delta$, described as a function of $d$, is obtained as in equation 22, where $t_{PFA}$ is the thickness value obtained from PFA with perturbation parameter $\epsilon=0.1$, and $t_{PFA+NTLO}$ is the thickness value received from the PFA method with NTLO correction.
@@
The fractional deviation value indicates that the PFA with NTLO provides a more realistic estimate that accounts for the curvature-dependent effects of the arc geometry than simply calculating the force with the PFA with a perturbation parameter $\epsilon=0.1$ The mean value of the fractional deviation results in $98.99\%$. That is, the thickness range, as shown in table 3, is $t_{PFA}\in[10^{-12};10^{-10}]$. It is important to note that the fractional deviation obtains such a high value due to the difference in the scale, in addition to the unphysical quantities which the PFA with a perturbation parameter sets to $\epsilon=0.1$. Detailed data for all values of the distance range $d\in[0.1-1]\mu m$ are shown in Fig. 4(a) and Fig. 4(b). However, it should be noted that violating the PFA condition $d<<R$ can result in significantly higher fractional deviations.
@@ 4.1 Applications
Although the Casimir pull-in is a disadvantage of the Casimir effect, the Casimir force has practical applications in MEMS, as it is a source of energy at the micro- and sub-micron levels. Attempts have been made to tune the Casimir force to obtain control over its magnitude, with applications such as shifting the force from attractive to repulsive. In addition to the method involving tuning the Casimir force by dielectric control as observed in detail in Section 1.3 of this paper ([16], [17]), [32] reviews tuning the Casimir force by external magnetic fields. This suggests that the restoring force can be obtained by using tuning methods. Enabling the use of tuning as a restoration mechanism creates the possibility of incorporating cyclic use in MEMS where the Casimir force performs work, and a restoring force brings the system to the initial position. Such cyclic interactions have applications in casimir-driven actuators. Designing a bridge-type MEMS switch with the required thickness and active Casimir force tuning would enable the use of the Casimir effect as an energy source. This is an alternative to cantilever switch design, which uses quantum energy and overcomes stiction. These switches find their best applications in vacuum because the Casimir force is the strongest. The required environment enables their use in MEMS for space applications. Harnessing the Casimir effect as an energy source allows for a lower power input; consequently, the system could operate longer for the equivalent power input. As reported by [33], MEMS switch solid-state equivalents require higher drive power.
Curvature reversal allows the design of a micropump. The micropump is composed of the arc-plate geometry, where the Casimir force drives the motion, making it a non-mechanical pump, as it uses Casimir energy to produce kinetic energy. Returning to the initial position of curvature, that is, reversing the curvature of the nanomembrane from convex to concave with respect to the plate can be achieved using the proposed restoring forces acquired from Casimir force tuning. Micropumps have many applications, some of which include cooling in nanoelectronics and fuel delivery in nanorobotics. Moreover, micropumps have applications in the biomedical sphere for drug delivery, as in [34]. All three may benefit from the incorporation of the Casimir effect as either an energy or pressure source.
@@ Conclusions
The Casimir effect was observed for the arc-plate geometry. Specifically, a theoretical estimate of the Casimir energy required to bend a nanomembrane from the initial position, which is concave to the plate, to the final position, which is convex to the plate, was conducted. The Casimir energy is calculated using the PFA method with NTLO correction for an array of 1000 distances in the range $d\in[0.1-1]\mu m$; where the radius of the arc is $100\mu m$ in order to satisfy the PFA condition, i.e. $d<<R$. The maximum thickness was computed by comparing the Casimir energy to the estimated bending energy of the nanomembranes for gold and silver. The maximum thickness to be able to perform curvature reversal is obtained in the range $9.636$ nm $7.817\times 10^{-1}$ nm for silver and $9.532$ nm $7.732\times 10^{-1}$ nm for gold. The precision of the NTLO correction was evaluated in comparison with PFA with a perturbation parameter $\epsilon=0.1$. Applications in MEMS are discussed, including methods for overcoming stiction in MEMS cantilever switches and Casimir-driven devices such as micropumps. Future work will focus on coupling the present analytical framework with finite-element simulations to investigate Casimir-induced deformation and stability under realistic boundary conditions.

### v2

@@ 2.4 Procedure
In classical plate and shell theory, curvature reversal corresponds to a buckling-type instability whose critical condition depends on boundary constraints and in-plane stresses. A full buckling analysis under Casimir loading would require solving nonlinear equilibrium equations and identifying bifurcation points for curved beams or shells, typically involving eigenvalue problems and numerical methods (emam_review_2022, ). Such an analysis is beyond the scope of our study and is left for future investigation.
@@ 2.4.1 Calculating Casimir Energy
Fosco’s correction was implemented by applying a derivative expansion to the first two terms as described in equation 13 where $U_{DE}$ is the Casimir energy including the NTLO correction, $\hbar$ is the reduced Planck constant, $c$ is the speed of light in vacuum, $x_{||}$ denotes the coordinates parallel to the reference surface over which the integration is performed, $\psi(d)$ is the local separation, $d$, between the curved surface and the opposing plate (equation 14), and $(\partial\psi)^{2}$ represents the sum of the squared partial derivatives of $\psi$ with respect to the parallel coordinates, i.e., $(\partial\psi)^{2}=(\partial\psi/\partial x)^{2}+(\partial\psi/\partial y)^{2}$ (fosco_proximity_2011, ).
The geometric function describes the arc, as expressed in equation 14. Restrictions were imposed, namely [3<y<3][-3<y<3], to prevent contact between the two fixed surfaces. As described in fosco_proximity_2011 , the first term in the derivative expansion corresponds to PFA. Thus, if one approximates the second term as a small perturbation parameter, $\epsilon$, by perturbation theory, $U_{DE}$ reduces the expression to the PFA.
NTLO corrections serve to better approximate the Casimir energy for the arc geometry due to its geometrical properties, accounting for variations in the local slope where distances change nonlinearly. These corrections are derived using a derivative expansion, which is valid when the local surface curvature varies slowly compared to the separation distance. This condition is satisfied in our study for all distances $d\in[0.1,1]\,\mu\mathrm{m}$ with a curvature radius $R=100\,\mu\mathrm{m}$. This ensures that the NTLO correction remains controlled and physically meaningful. Experimental verification of the accuracy of PFA with NTLO corrections is left for future work.
@@ 2.4.2 Calculating Bending Energy
The bending energy, $U_{bending}$, is obtained using equation 15, where $u$ is the bending strain energy and $A$ is the area (wierzbicki_tomasz_2081j16230j_nodate, ). Because we obtain a numerical value for $u$, the equation simplifies to equation 16, where $A$ simplifies to $L$, the length of the arc. To evaluate the arc length, $L$ is defined in equation 17, where $y_{min}=-3\mu m$ and $y_{max}=3\mu m$. The length of the arc was $L=6\mu m$.
@@
Table 1 shows the values of $E$, Young’s modulus, and $\upsilon$, Poisson’s ratio, for thin gold and thin silver films (noel_review_2016, ), (ruud_elastic_1991, ).
@@ 2.5 Algorithm for Critical Thickness Evaluation
The procedure for evaluating the critical thickness can be summarized in the following algorithm.
@@
7:Apply derivative expansion correction fosco_proximity_2011 :
@@ 3.1 Estimating the maximum thickness using PFA with NTLO corrections
Including higher-order corrections to the PFA contributes to a more realistic estimate of the results. However, our analysis shows that the higher-order correction is negligible, modifying the results only at the level of the fourth significant digit.
@@
We quantify this estimate by the fractional deviation $\delta$, described as a function of $d$, which is obtained as in equation 22, where $t_{PFA}$ is the thickness value obtained from PFA with perturbation parameter $\epsilon=0.1$, and $t_{PFA+NTLO}$ is the thickness value received from the PFA method with NTLO correction.
@@
Different geometries could result in a higher fractional deviation value that accounts for curvature-dependent effects. Regarding the arc-to-plate geometry in this study and its quasi-parallel nature, the PFA with perturbation parameter $\epsilon=0.1$ serves as a good approximation, while PFA with NTLO further precises it. However, we consider the NTLO correction effects as negligible. The thickness range of the PFA perturbation parameter $\epsilon=0.1$ can be observed in table 3. Detailed data for all values of the distance range $d\in[0.1-1]\mu m$ are shown in Fig. 4(a) and Fig. 4(b). Since the deviation can be expressed as a ratio of energies that factors out the material-dependent constants, it is identical for both silver and gold. However, it should be noted that violating the PFA condition $d<<R$ can result in significantly higher fractional deviations.
@@ 4.1 Applications
Although the Casimir pull-in is a disadvantage of the Casimir effect, the Casimir force has practical applications in MEMS, as it is a source of energy at the micro- and sub-micron levels. Attempts have been made to tune the Casimir force to obtain control over its magnitude, with applications such as shifting the force from attractive to repulsive. In addition to the method involving tuning the Casimir force by dielectric control as observed in detail in Section 1.3 of this paper ((munday_measured_2009, ), (zhao_stable_2019, )), (passante_quantum_2025, ) reviews tuning the Casimir force by external magnetic fields. This suggests that the restoring force can be obtained by using tuning methods. Enabling the use of tuning as a restoration mechanism creates the possibility of incorporating cyclic use in MEMS where the Casimir force performs work, and a restoring force brings the system to the initial position. Such cyclic interactions have applications in casimir-driven actuators. Designing a bridge-type MEMS switch with the required thickness and active Casimir force tuning would enable the use of the Casimir effect as an energy source. This is an alternative to cantilever switch design, which uses quantum energy and overcomes stiction. These switches find their best applications in vacuum because the Casimir force is the strongest. The required environment enables their use in MEMS for space applications. Harnessing the Casimir effect as an energy source allows for a lower power input; consequently, the system could operate longer for the equivalent power input. As reported by (pacheco_design_2000, ), MEMS switch solid-state equivalents require higher drive power.
Curvature reversal allows the design of a micropump. The micropump is composed of the arc-plate geometry, where the Casimir force drives the motion, making it a non-mechanical pump, as it uses Casimir energy to produce kinetic energy. Returning to the initial position of curvature, that is, reversing the curvature of the nanomembrane from convex to concave with respect to the plate can be achieved using the proposed restoring forces acquired from Casimir force tuning. Micropumps have many applications, some of which include cooling in nanoelectronics and fuel delivery in nanorobotics. Moreover, micropumps have applications in the biomedical sphere for drug delivery, as in (lo_passive_2009, ). All three may benefit from the incorporation of the Casimir effect as either an energy or pressure source.
@@ Conclusions
The Casimir effect was observed for the arc-plate geometry. Specifically, a theoretical estimate of the Casimir energy required to bend a nanomembrane from the initial position, which is concave to the plate, to the final position, which is convex to the plate, was conducted. The Casimir energy is calculated using the PFA method with NTLO correction for an array of 1000 distances in the range $d\in[0.1-1]\mu m$; where the radius of the arc is $100\mu m$ in order to satisfy the PFA condition, i.e. $d<<R$. The maximum thickness was computed by comparing the Casimir energy to the estimated bending energy of the nanomembranes for gold and silver. The maximum thickness to be able to perform curvature reversal is obtained in the range $9.636$ nm $7.817\times 10^{-1}$ nm for silver and $9.532$ nm $7.732\times 10^{-1}$ nm for gold. The precision of the NTLO correction was evaluated in comparison with PFA with a perturbation parameter $\epsilon=0.1$. Applications in MEMS are discussed, including methods for overcoming stiction in MEMS cantilever switches and Casimir-driven devices such as micropumps. Future work could focus on coupling the present analytical framework with finite-element simulations to investigate Casimir-induced deformation and stability under realistic boundary conditions. This work provides a theoretical framework for Casimir forces in nanofilms, which could serve as a basis for future numerical simulations and experimental studies.
@@ Acknowledgements
Funding
This work was funded by the Vicerrectorado de Investigación (VRI) at the Pontificia Universidad Católica del Perú through the Dirección de Fomento de la Investigación (Project ID: PI1287).